OpenStudy (anonymous):

Can someone help me understand PEMDAS combined with Distributive property? My teacher keeps saying Im making the same mistake: Combining what cant be combined. Help?

OpenStudy (anonymous):

Do you know what pemdas means?

OpenStudy (anonymous):

Parenthases, exponents, multiplication, division, addition, subtraction. AKA Please Excuse My Dear Aunt Mary

hero (hero):

Show An example of your mistake

OpenStudy (anonymous):

ok. 2g-6-8g=9-g+5 my answer: 2g-6-8g=9-g+5 10g - -6=9g+5 -10 -10 g-6=1+5 -6g = 6 6 6 =1 I dont get what im doing wrong?

OpenStudy (jackthegreatest):

u did the g's wrong

OpenStudy (jackthegreatest):

2g-8g is not 10g

OpenStudy (jackthegreatest):

and 9-g is not 9g

OpenStudy (jackthegreatest):

2g-6-8g=9-g+5 -6g-6 =-g+14 -20=5g g=-4

OpenStudy (jackthegreatest):

@magikgirl18 i hope this helps : )

OpenStudy (anonymous):

oh.....I added instead of subtracting, and screwed up my gs! Ok, thank you!

OpenStudy (jackthegreatest):

welcome! no worries! just remember ur mistakes so u dont mess up on tests :)

OpenStudy (anonymous):

ok!

OpenStudy (jackthegreatest):

: )